A, B and C are bidding for a contract. It is believed that A has exactly half the chance that B has, B in turn 4/5th as likely as C to win the contract. What is the probability for each to win the contract?


1
Expert's answer
2021-02-26T00:54:02-0500

Let probability of C winning be x

Therefore probability of B winning is "\dfrac{4x}{5}"

and probability of A winning is "\dfrac{4x}{10}"


Sum of the probabilities should be equal to 1

So, "\dfrac{4x}{5} +\dfrac{4x}{10}+x=1"

"\Rightarrow x=\dfrac{5}{11}"


Therefore,

A has a probability of "\dfrac{2}{11}" for winning the contract.


B has a probability of "\dfrac{4}{11}" for winning the contract.


C has a probability of "\dfrac{5}{11}" for winning the contract.